OpenStudy (anonymous):

Tommy runs 3/4 times as fast as Johnny. Also, Tommy gives Johnny a lead of 77 metre. Both of them reach at the same time. Find the length of the track.

OpenStudy (anonymous):

sorry its 3*3/4

OpenStudy (anonymous):

77=15/4x, 308/15=x

OpenStudy (anonymous):

15/4*x=x+77

OpenStudy (anonymous):

15x=4x+308 x=28

OpenStudy (anonymous):

so 105 meter